# Multimodal Recurrent Ensembles for Predicting Brain Responses to Naturalistic Movies (Algonauts 2025)

### v3

@@ 1 Introduction
Understanding how complex, naturalistic stimuli drive human brain activity is a core question in cognitive computational neuroscience naselaris_encoding_2011 ; yamins_performance-optimized_2014 ; kriegeskorte_deep_2015 ; doerig_neuroconnectionist_2023 . Recent advances in deep learning and systems neuroscience have produced models that map rich stimulus features onto distributed cortical responses. The Algonauts Project 2025 aims at supporting progress in this area by providing a large open dataset and a clear community benchmark for predicting fMRI BOLD signals boyle_courtois_2023 ; gifford_algonauts_2025 from naturalistic movie stimuli. With a common dataset, standardized metrics and a collaborative challenge format, the project allows researchers to compare ideas head-to-head and drives methods development towards more accurate, biologically grounded accounts of how the brain interprets the real world.
@@ 2 Related Work
Recent advances in computational neuroscience have shown that deep learning models provide powerful tools for predicting brain responses to multimodal, naturalistic stimuli naselaris_encoding_2011 ; yamins_performance-optimized_2014 . While initial studies primarily encoded fMRI responses to isolated sensory inputs such as static images or speech kriegeskorte_deep_2015 ; wen_neural_2018 , recent research highlights the benefits of integrating vision, audio, and language into unified brain-encoding models gifford_algonauts_2025 . Models combining audiovisual features have notably improved prediction accuracy in higher-order cortical regions cichy_algonauts_2021 .
Capturing temporal dynamics is essential since natural stimuli induce brain activity evolving over extended timescales. Recurrent neural networks (RNNs), particularly long short-term memory (LSTM) hochreiter_long_1997 and gated recurrent units (GRUs) chung_empirical_2014 , have successfully modeled temporal dependencies in brain data. Transformer-based architectures, capable of encoding longer temporal contexts, have also gained attention, though RNN-based methods remain prevalent due to their effectiveness and simplicity adeli_transformer-based_2023 .
Benchmarking initiatives such as the Algonauts Project significantly accelerated progress through standardized datasets and evaluation methods gifford_algonauts_2025 . Participants frequently adopt ensemble learning—averaging predictions from multiple diverse models—to enhance accuracy and robustness nguyen_multiobjective_2023 . Curriculum learning, progressively shifting training focus from simpler to more complex brain regions, has similarly improved model stability and predictive power bengio_curriculum_2009 . Our approach integrates these strategies to provide an effective multimodal baseline for future brain-encoding benchmarks.
@@ 4.1.1 Multimodal Feature Extraction
To obtain rich representations of each stimulus modality, we leverage state-of-the-art pretrained models. For the visual input, we extract features from each video using four complementary encoders—SlowFast feichtenhofer_slowfast_2019 , VideoMAE tong_videomae_2022 , Swin Transformer liu2021swintransformerhierarchicalvision , and CLIP radford2021learningtransferablevisualmodels —each trained on large-scale video or image datasets. Visual features are computed on 1.49‑second clips and then time‑aligned to the fMRI time points. For the auditory stream, we employ multiple pretrained audio models—HuBERT hsu_hubert_2021 and WavLM chen_wavlm_2022 for self‑supervised speech/audio representations, and CLAP elizalde_clap_2023 for semantic audio embeddings—also extracted from the same 1.49‑second windows and aligned to the fMRI. The text input is represented using two language models applied to dialogue transcripts: a base BERT devlin_bert_2019 model for local semantic features (by feeding the last $n$ tokens based on the maximum token length of the language model) and a Longformer beltagy_longformer_2020 for longer‑range context. To ensure continuity of the transcripts of each episode and boost the performance of the first time stamps of the segments (e.g. episode split), we prepend the previous episode’s transcripts (if available) when extracting language features.
@@ 4.4 Validation
One training heuristic we explored was inspired by the hierarchical processing in the brain. Primary sensory areas encode basic stimulus features, whereas higher-order regions accumulate and integrate information over much longer timescales. As a result, low‑level sensory signals tend to be more predictable, while forecasting the responses of higher‑order areas is more challenging. We hypothesised that emphasising early sensory ROIs at the start of training could guide the model to first learn low-level stimulus-response mappings, before focusing on more abstract regions. To implement this, we devised a dynamic loss weighting scheme that gradually shifts focus from early to late-processing ROIs over training epochs. Concretely, we predefined a set of “early-processing” parcels (covering Visual and Somatomotor Network from Schaefer 2018 parcellation schaefer_2018 ) based on the provided ROI labels. During training, we split the loss into two components: $\mathcal{L}_{E}$ for the early-ROI subset and $\mathcal{L}_{L}$ for the remaining late-ROI set. We then applied a time-varying weight: at the beginning, the loss weight $w_{E}$ for $\mathcal{L}_{E}$ is higher (0.55) and $w_{L}=1-w_{E}$ is lower (0.45), biasing training towards fitting early ROIs. As epochs progress, we linearly anneal these weights towards a balanced emphasis (e.g. $w_{E}$ down to 0.5 and $w_{L}$ up to 0.5). By the end of training, late ROIs receive equal or greater attention. In our experiments, this strategy improved the model’s convergence on challenging ROIs and yielded a small boost in overall correlation performance (especially for mid-level and higher cortical areas) compared to a uniform loss weighting.

### v4

@@ 1 Introduction
Understanding how complex, naturalistic stimuli drive human brain activity is a core question in cognitive computational neuroscience (Naselaris et al., 2011; Yamins and DiCarlo, 2016; Kriegeskorte, 2015; Doerig et al., 2023). Recent advances in deep learning and systems neuroscience have produced models that map rich stimulus features onto distributed cortical responses. The Algonauts Project 2025 aims at supporting progress in this area by providing a large open dataset and a clear community benchmark for predicting fMRI BOLD signals (Boyle et al., 2023; Gifford et al., 2025) from naturalistic movie stimuli. With a common dataset, standardized metrics and a collaborative challenge format, the project allows researchers to compare ideas head-to-head and drives methods development towards more accurate, biologically grounded accounts of how the brain interprets the real world.
@@ 2 Related Work
Recent advances in computational neuroscience have shown that deep learning models provide powerful tools for predicting brain responses to multimodal, naturalistic stimuli (Naselaris et al., 2011; Yamins and DiCarlo, 2016). While initial studies primarily encoded fMRI responses to isolated sensory inputs such as static images or speech (Kriegeskorte, 2015; Wen et al., 2017), recent research highlights the benefits of integrating vision, audio, and language into unified brain-encoding models (Gifford et al., 2025). Models combining audiovisual features have notably improved prediction accuracy in higher-order cortical regions (Cichy et al., 2019).
Capturing temporal dynamics is essential since natural stimuli induce brain activity evolving over extended timescales. Recurrent neural networks (RNNs), particularly long short-term memory (LSTM) (Hochreiter and Schmidhuber, 1997) and gated recurrent units (GRUs) (Chung et al., 2014), have successfully modeled temporal dependencies in brain data. Transformer-based architectures, capable of encoding longer temporal contexts, have also gained attention, though RNN-based methods remain prevalent due to their effectiveness and simplicity (Adeli et al., 2023).
Benchmarking initiatives such as the Algonauts Project significantly accelerated progress through standardized datasets and evaluation methods (Gifford et al., 2025). Participants frequently adopt ensemble learning—averaging predictions from multiple diverse models—to enhance accuracy and robustness (Nguyen et al., 2023). Curriculum learning, progressively shifting training focus from simpler to more complex brain regions, has similarly improved model stability and predictive power (Bengio et al., 2009). Our approach integrates these strategies to provide an effective multimodal baseline for future brain-encoding benchmarks.
@@ 4.1.1 Multimodal Feature Extraction
To obtain rich representations of each stimulus modality, we leverage state-of-the-art pretrained models. For the visual input, we extract features from each video using four complementary encoders—SlowFast (Feichtenhofer et al., 2019), VideoMAE (Tong et al., 2022), Swin Transformer (Liu et al., 2021), and CLIP (Radford et al., 2021)—each trained on large-scale video or image datasets. Visual features are computed on 1.49‑second clips and then time‑aligned to the fMRI time points. For the auditory stream, we employ multiple pretrained audio models—HuBERT (Hsu et al., 2021) and WavLM (Chen et al., 2022) for self‑supervised speech/audio representations, and CLAP (Elizalde et al., 2022) for semantic audio embeddings—also extracted from the same 1.49‑second windows and aligned to the fMRI. The text input is represented using two language models applied to dialogue transcripts: a base BERT (Devlin et al., 2019) model for local semantic features (by feeding the last $n$ tokens based on the maximum token length of the language model) and a Longformer (Beltagy et al., 2020) for longer‑range context. To ensure continuity of the transcripts of each episode and boost the performance of the first time stamps of the segments (e.g. episode split), we prepend the previous episode’s transcripts (if available) when extracting language features.
@@ 4.4 Validation
One training heuristic we explored was inspired by the hierarchical processing in the brain. Primary sensory areas encode basic stimulus features, whereas higher-order regions accumulate and integrate information over much longer timescales. As a result, low‑level sensory signals tend to be more predictable, while forecasting the responses of higher‑order areas is more challenging. We hypothesised that emphasising early sensory ROIs at the start of training could guide the model to first learn low-level stimulus-response mappings, before focusing on more abstract regions. To implement this, we devised a dynamic loss weighting scheme that gradually shifts focus from early to late-processing ROIs over training epochs. Concretely, we predefined a set of “early-processing” parcels (covering Visual and Somatomotor Network from Schaefer 2018 parcellation (Schaefer et al., 2018)) based on the provided ROI labels. During training, we split the loss into two components: $\mathcal{L}_{E}$ for the early-ROI subset and $\mathcal{L}_{L}$ for the remaining late-ROI set. We then applied a time-varying weight: at the beginning, the loss weight $w_{E}$ for $\mathcal{L}_{E}$ is higher (0.55) and $w_{L}=1-w_{E}$ is lower (0.45), biasing training towards fitting early ROIs. As epochs progress, we linearly anneal these weights towards a balanced emphasis (e.g. $w_{E}$ down to 0.5 and $w_{L}$ up to 0.5). By the end of training, late ROIs receive equal or greater attention. In our experiments, this strategy improved the model’s convergence on challenging ROIs and yielded a small boost in overall correlation performance (especially for mid-level and higher cortical areas) compared to a uniform loss weighting.